OpenStudy (anonymous):

Can someone walk me through this problem? The area of a rectangular pool is given by the trinomial 2y^2+y-3. whats the demensions of the pool?

OpenStudy (mertsj):

Are you sure there is not a typo in your post?

OpenStudy (anonymous):

sorry i mixed up 2 questions

Directrix (directrix):

We need to factor this: 2y^2+y-3

Directrix (directrix):

Two number that multiply to -6 and add to 1

Directrix (directrix):

2 y^2 + 1*y - 3 = 2 y^2 + 3y - 2y - 3

OpenStudy (anonymous):

my sister woke up. i wont be on until later 2 finish this

Directrix (directrix):

Okay. Catch you later.

Directrix (directrix):

2 y^2 + 1*y - 3 = 2 y^2 + 3y - 2y - 3 = 2 y^2 - 2y + 3y - 3 = 2y(y - 1) + 3(y -1) = (y - 1)* (2y + 3) --> pool dimensions @JaneDoe100

OpenStudy (anonymous):

Thanks! Really appreciate all the time you put into helping me...